OpenStudy (gabby1214):

solve 3/7x=6 The solution is x=?(Integer or simplified fraction)

pooja195 (pooja195):

Multiply by the reciprocal \[\huge~\rm~ \frac{ 7 }{ 3 }*\frac{ 3 }{ 7x }=\frac{ 7 }{ 3 }*6\]